You want to have a million dollars in the bank when you retire. You think you can save $5 000 this year, and increase that by $100 every subsequent year, in a bank that offers you 5% interest. If you make your first deposit in a year's time, how many years will it be from now before you can retire?

Select one:

a. 45

b. 50

c. 35

d. 40

e. 30
Answer #1

Let deposit is required for n yrs, then future worth of deposit

F = 5000 * (F/A, 5%,n) + 100 *(F/G, 5%, n)

As per the given condition this is to be equal to 1000000

5000 * (F/A, 5%,n) + 100 *(F/G, 5%, n) = 1000000

Using trail and error method

When n = 35, value of 5000 * (F/A, 5%,n) + 100 *(F/G, 5%, n) = 562242.15

When n = 45, value of 5000 * (F/A, 5%,n) + 100 *(F/G, 5%, n) = 1027901.09

This is just above 1m, as question does not required it to solve uptill this step only

So correct answer is option A. 45